The height of a triangle is one-half the length of the base. The area of the triangle 144in exponent 2. Find the height of the triangle.
Answer by stanbon(75887)

Area of a triangle = (1/2) (base)(height)
Let the base be "b"; then the height is (1/2)b
Equation:
144 sq in = (1/2)(b)((1/2)(b))
144 = (1/4)b^2
(4)(144) = b^2
b = 2(12)
b = 24 inches
h = 12 inches